Circular No. 6749

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dg IN ANONYMOUS GALAXY Q.-y. Qiao, Y.-l. Qiu, and J.-y. Hu, Beijing Astronomical Observatory

(BAO); A. Esamdin, Urmqi Astronomical Station; and Y. Zhang, Beijing Normal University, on behalf of the BAO Supernova Survey, report their discovery of an apparent supernova found on unfiltered CCD images taken on Sept. 27.72 (mag ∼ 16.7) and 28.60 UT (mag 16.4) with the BAO 0.60-m reflector. SN 1997dg is located near the center of the host galaxy, at α = 23^{h}40^{m}14^{s}.21, δ = +26^{o}12′11′′.8 (equinox 2000.0). Images of the same field taken on Sept. 4 (limiting mag ∼ 19.0) show no star at this position.

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Center for Astrophysics, report that a spectrum of SN 1997dg taken by J. Luu on Sept. 30.3 UT at the Multiple Mirror Telescope reveals that it is a type-Ia supernova near maximum light. The approximate redshift is 0.03, derived from the strong Si II 615-nm feature. Images taken by N. Grogin on Sept. 29.4 at the Whipple Observatory 1.2-m telescope yield magnitudes B = 17.2, R = 16.9 (measured relative to nearby USNO-A1.0 star 1125.20187894).

V2116 OPHIUCHI M. G. Pereira, F. Jablonski, and J. Braga, Instituto Nacional de Pes-

quisas Espaciais; and D. Chakrabarty, Massachusetts Institute of Technology, write: “Following the detection of optical pulsations close to the rotational period of the neutron star in V2116 Oph = GX 1+4 (IAUC 6489; Jablonski et al. 1997, Ap.J. 482, L171), we made additional optical-period measurements during 1997 with a 0.6-m telescope (+ CCD photometer; 20–30-s time resolution in the R band). Fourteen different determinations yield a barycentric period ranging from 126.36 ± 0.19 s on May 18.25 UT to 127.50 ± 0.16 s on Aug. 20.08. Although V2116 Oph was in a bright optical state (R = 15.5–15.9) during most of July, no optical pulsations were detected in the interval July 24–29, to a detection limit of ∼ 0.3 percent in modulation amplitude. The optical measurements for the interval May– Aug. indicate a spin-down rate of +4.0 ± 0.6 s/yr. Preliminary analysis of 20–60-keV x-ray timing data from the CGRO/BATSE all-sky monitor shows that the x-ray pulse period of GX 1+4 was 126.438 ± 0.008 s on May 18.5 and 126.986 ± 0.008 s on July 11.5. The pulsar was rapidly spinning down at a relatively steady rate of +3.9 s/yr from early May through early July, and then abruptly decreased its spin-down rate to +2.0 s/yr in early July. At present, it continues to spin down at this rate. The May–July rate is consistent with the value indicated by the optical measurements.”

1997 September 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green